Explain to a grade 4 class why pencils on the table are 3/10 of the pencils, not 3/7


1
Expert's answer
2021-05-28T09:58:26-0400

When we have 3 pencils on the table and rest of the 7 pencils are on the floor.


Then the total number of pencils "=" pencils on table "+" pencils on the floor

total number of pencils "=3+7=10"


Now Factor of pencils on the table = pencils on the table divided by total pencils

According to this:


Factor of pencils on the table "= \dfrac{3}{10}"( of the total pencils)


OR


In language we can say that only 3 pencils out of total 10 pencils are on the table ,so


there is "\dfrac{3}{10}" parts of the total pencils present on the table.


In above explanation there is no mention of "\dfrac{3}{7}" so we can say that the answer can not be "\dfrac{3}{7}" .as denominator always contains the total no of pencils present(that is 10 and not 7).